OpenStudy (anonymous):

Solve. 64^x = 16^x–1

OpenStudy (king):

x=5/2?

OpenStudy (anonymous):

use \[64=4^3,16=4^2\] to write both sides with the same base

OpenStudy (anonymous):

64^x = 16^x–1